# fold/.style=folder, grow’=0, s sep=0.1pt, arr/.style=edge=-latex

### v3

@@ VII Outlook
In our presentation of our RL framework, we focused on the exact synthesis of unitaries using the Clifford+T gate set. With some further improvements, the scope of our agent’s application can be expanded. Firstly, our approach does not preserve the MCTS tree following the execution of actions, meaning the search tree is rebuilt from scratch at every iteration. We aim to address this limitation in our future work. Moreover, akin to QCOpt, augmenting the action masking process to encompass additional circuit equivalence relations might be feasible. A further enhancement could involve adopting a more advanced neural network design, like a transformer network, which could consider the circuit’s configuration in addition to the target unitary, potentially elevating the agent’s efficacy. An intriguing inquiry is whether a more effective method exists for representing the unitary as input to the neural network.
Transitioning to approximate synthesis theoretically entails dropping the constraint that $C_{HS}$ be precisely zero for two unitaries. Another consideration is to incorporate unitaries that utilize ancilla qubits. This is because certain unitaries, like the controlled-T, may necessitate an ancilla qubit for their synthesis. Additionally, using an ancilla qubit could reduce the depth of the synthesized circuit.
@@ Acknowledgement
We are very grateful to P.  Mukhopadhyay for helpful discussions and for sharing the code with us, which was used to generate the results for the MIN-T-SYNTH method shown in Table I. We would also like to thank V. Gheorghiu for sharing the Github repository with us for the approximate circuit synthesis in Clifford+T gates. We thank N. Quetschlich for interesting discussions related to quantum compilation using reinforcement learning. Finally, we would like to thank Q. Göttl for comments on the manuscript.

### v4

@@ VII Outlook
In our presentation of our RL framework, we focused on the exact synthesis of unitaries using the Clifford+T gate set. With some further improvements, the scope of our agent’s application can be expanded. Firstly, our approach does not preserve the MCTS tree following the execution of actions, meaning the search tree is rebuilt from scratch at every iteration. We aim to address this limitation in our future work. A further enhancement could involve adopting a more advanced neural network design, like a transformer network, which could consider the circuit’s configuration in addition to the target unitary, potentially elevating the agent’s efficacy. An intriguing inquiry is whether a more effective method exists for representing the unitary as input to the neural network. We would also like to alleviate a limitation of our benchmarking results by comparing other RL based methods for synthesizing unitaries such as Ref. [57].
Transitioning to approximate synthesis theoretically entails dropping the constraint that $C_{HS}$ be precisely zero for two unitaries, which will be the next obvious step. Another consideration is to incorporate unitaries that utilize ancilla qubits, which could potentially reduce the depth of the synthesized circuit.
@@ Acknowledgement
We are very grateful to P.  Mukhopadhyay for helpful discussions and for sharing the code with us, which was used to generate the results for the MIN-T-SYNTH method shown in Table II. We would also like to thank V. Gheorghiu for sharing the Github repository with us for the approximate circuit synthesis in Clifford+T gates. We thank N. Quetschlich for interesting discussions related to quantum compilation using reinforcement learning. Finally, we would like to thank Q. Göttl for comments on the manuscript.